OpenStudy (anonymous):

how do i find the y intercept when I have 2 points: (2,-3) and (-4,1)?

OpenStudy (anonymous):

first u have to find the slope then write the equation and solve for y..

OpenStudy (anonymous):

then to find the y-intercept u have to set x=0

OpenStudy (callisto):

Set up an equation using y-y1 y2-y1 ---- = ------ x-x1 x2 - x1 Put (x1,y1) = (2,-3) and (x2,y2) =(-4,1) into the equation What can you get so far?

OpenStudy (anonymous):

i have the slope of -2/3

OpenStudy (anonymous):

yea thts right!

OpenStudy (anonymous):

now make an equation u can pick points from any one of the bracket

OpenStudy (anonymous):

(y+3)=-2/3(x-2) <- like this, now solve for y and set x=0 to get the y intercept

OpenStudy (anonymous):

where did y+3 come from? shouldnt it be y-3?

OpenStudy (anonymous):

coz if u see in the question the points are (2,-3) here - times -3 is +3

OpenStudy (anonymous):

u know the eq right? (y1-y2)=m(x1-x2) so (y-(-3)) => (y+3)

OpenStudy (anonymous):

oh, ok, i always get confused with the signs:)